% Generated by IEEEtran.bst, version: 1.14 (2015/08/26)
\begin{thebibliography}{1}
\providecommand{\url}[1]{#1}
\csname url@samestyle\endcsname
\providecommand{\newblock}{\relax}
\providecommand{\bibinfo}[2]{#2}
\providecommand{\BIBentrySTDinterwordspacing}{\spaceskip=0pt\relax}
\providecommand{\BIBentryALTinterwordstretchfactor}{4}
\providecommand{\BIBentryALTinterwordspacing}{\spaceskip=\fontdimen2\font plus
\BIBentryALTinterwordstretchfactor\fontdimen3\font minus \fontdimen4\font\relax}
\providecommand{\BIBforeignlanguage}[2]{{%
\expandafter\ifx\csname l@#1\endcsname\relax
\typeout{** WARNING: IEEEtran.bst: No hyphenation pattern has been}%
\typeout{** loaded for the language `#1'. Using the pattern for}%
\typeout{** the default language instead.}%
\else
\language=\csname l@#1\endcsname
\fi
#2}}
\providecommand{\BIBdecl}{\relax}
\BIBdecl

\bibitem{neng-chao_research_2018}
\BIBentryALTinterwordspacing
L.~Y.~U. Neng-chao, C.~A.~O. Yue, Q.~I.~N. Ling, and W.~U. Chao-zhong, ``Research on the {Effectiveness} of {Driving} {Workload} {Based} on {Traffic} {Sign} {Information} {Volume},'' \emph{China Journal of Highway and Transport}, vol.~31, no.~8, pp. 165--172, 2018, publisher: China Journal of Highway and Transport. [Online]. Available: \url{https://zgglxb.chd.edu.cn/EN/Y2018/V31/I8/165}
\BIBentrySTDinterwordspacing

\bibitem{yiu_keeping_2025}
\BIBentryALTinterwordspacing
C.~Y. Yiu, K.~K.~H. Ng, Q.~Li, and X.~Yuan, ``\BIBforeignlanguage{en}{Keeping {Pilots} in the {Loop}: {An} {Explainable} {Spatiotemporal} {EEG}-{Driven} {Deep} {Learning} {Framework} for {Adaptive} {Automation} in {Cruising} {Flight} {Phase}},'' \emph{\BIBforeignlanguage{en}{IEEE Transactions on Intelligent Transportation Systems}}, pp. 1--14, 2025. [Online]. Available: \url{https://ieeexplore.ieee.org/document/11008888/}
\BIBentrySTDinterwordspacing

\bibitem{lin_hatnet_2025}
\BIBentryALTinterwordspacing
S.~Lin, C.~Fan, D.~Han, Z.~Jia, Y.~Peng, and S.~Kwong, ``\BIBforeignlanguage{en}{{HATNet}: {EEG}-{Based} {Hybrid} {Attention} {Transfer} {Learning} {Network} for {Train} {Driver} {State} {Detection}},'' \emph{\BIBforeignlanguage{en}{IEEE Transactions on Cybernetics}}, vol.~55, no.~5, pp. 2437--2450, May 2025. [Online]. Available: \url{https://ieeexplore.ieee.org/document/10908421/}
\BIBentrySTDinterwordspacing

\bibitem{weng_self-supervised_2025}
\BIBentryALTinterwordspacing
W.~Weng, Y.~Gu, S.~Guo, Y.~Ma, Z.~Yang, Y.~Liu, and Y.~Chen, ``\BIBforeignlanguage{en}{Self-supervised {Learning} for {Electroencephalogram}: {A} {Systematic} {Survey}},'' \emph{\BIBforeignlanguage{en}{ACM Computing Surveys}}, p. 3736574, May 2025. [Online]. Available: \url{https://dl.acm.org/doi/10.1145/3736574}
\BIBentrySTDinterwordspacing

\bibitem{jiang2024large}
\BIBentryALTinterwordspacing
W.~Jiang, L.~Zhao, and B.~liang Lu, ``Large brain model for learning generic representations with tremendous {EEG} data in {BCI},'' in \emph{The Twelfth International Conference on Learning Representations}, 2024. [Online]. Available: \url{https://openreview.net/forum?id=QzTpTRVtrP}
\BIBentrySTDinterwordspacing

\bibitem{ma2024attention}
X.~Ma, W.~Chen, Z.~Pei, Y.~Zhang, and J.~Chen, ``Attention-based convolutional neural network with multi-modal temporal information fusion for motor imagery eeg decoding,'' \emph{Computers in Biology and Medicine}, p. 108504, 2024.

\bibitem{muse_cldrive}
P.~Angkan, B.~Behinaein, Z.~Mahmud, A.~Bhatti, D.~Rodenburg, P.~Hungler, and A.~Etemad, ``Multimodal brain–computer interface for in-vehicle driver cognitive load measurement: Dataset and baselines,'' \emph{IEEE Transactions on Intelligent Transportation Systems}, vol.~25, no.~6, pp. 5949--5964, 2024.

\bibitem{lawhernEEGNetCompactConvolutional2018}
V.~J. Lawhern, A.~J. Solon, N.~R. Waytowich, S.~M. Gordon, C.~P. Hung, and B.~J. Lance, ``{{EEGNet}}: A compact convolutional neural network for {{EEG-based}} brain--computer interfaces,'' \emph{Journal of Neural Engineering}, vol.~15, no.~5, p. 056013, Oct. 2018.

\bibitem{schirrmeisterDeepLearningConvolutional2018}
R.~T. Schirrmeister, J.~T. Springenberg, L.~D.~J. Fiederer, M.~Glasstetter, K.~Eggensperger, M.~Tangermann, F.~Hutter, W.~Burgard, and T.~Ball, ``Deep learning with convolutional neural networks for {{EEG}} decoding and visualization,'' \emph{Human Brain Mapping}, vol.~38, no.~11, pp. 5391--5420, Jun. 2018.

\end{thebibliography}

\documentclass[conference]{IEEEtran}
\IEEEoverridecommandlockouts
\usepackage{cite}
\usepackage{amsmath,amssymb,amsfonts}
\usepackage{algorithmic}
\usepackage{graphicx}
\usepackage{textcomp}
\usepackage{xcolor}
\usepackage{booktabs} 
\usepackage{siunitx}  
\usepackage{caption}  
\usepackage{bm}

\def\BibTeX{{\rm B\kern-.05em{\sc i\kern-.025em b}\kern-.08em
    T\kern-.1667em\lower.7ex\hbox{E}\kern-.125emX}}
\begin{document}

\title{Neuro-Informed Joint Learning Enhances Cognitive Workload Decoding in Portable BCIs}

\author{
\IEEEauthorblockN{Xiaoxiao Yang}
\IEEEauthorblockA{\textit{Peiyang Brain-Computer Interface \&} \\ 
\textit{Smart Health Inst.} \\ 
Xiamen, China \\ 
xiaoxiao.yang@ieee.org}
\and
\IEEEauthorblockN{Chao Feng}
\IEEEauthorblockA{\textit{Peiyang Brain-Computer Interface \&} \\
\textit{Smart Health Inst.} \\ 
Xiamen, China \\ 
charles\_feng@intretech.com}
\and
\IEEEauthorblockN{Jiancheng Chen}
\IEEEauthorblockA{\textit{Peiyang Brain-Computer Interface \&} \\
\textit{Smart Health Inst.} \\
Xiamen, China \\ 
xmcjc@intretech.com}
}
\maketitle

\begin{abstract}
Portable and wearable consumer-grade electroencephalography (EEG) devices, like Muse headbands, offer unprecedented mobility for daily brain-computer interface (BCI) applications, including cognitive load detection. However, the exacerbated non-stationarity in portable EEG signals constrains data fidelity and decoding accuracy, creating a fundamental trade-off between portability and performance. To mitigate such limitation, we propose MuseCogNet (Muse-based Cognitive Network), a unified joint learning framework integrating self-supervised and supervised training paradigms. In particular, we introduce an EEG-grounded self-supervised reconstruction loss based on average pooling to capture robust neurophysiological patterns, while cross-entropy loss refines task-specific cognitive discriminants. This joint learning framework resembles the bottom-up and top-down attention in humans, enabling MuseCogNet to significantly outperform state-of-the-art methods on a publicly available Muse dataset and establish an implementable pathway for neurocognitive monitoring in ecological settings.
\end{abstract}

\begin{IEEEkeywords}
BCI, EEG, Self-Supervised Learning, Cognitive Load, Wearable, Human Attention
\end{IEEEkeywords}

\section{Introduction}
Overloaded cognitive states in critical scenarios like driving are major contributors to car accidents~\cite{neng-chao_research_2018}, necessitating real-time cognitive workload detection. EEG-based BCIs offer promise due to non-invasiveness and temporal resolution. Yet researches based on medical-grade EEG devices~\cite{yiu_keeping_2025, lin_hatnet_2025} face deployment barriers: their complex setups require expertise and continuous monitoring. Consumer-grade devices like Muse headbands provide accessibility but introduce amplified non-stationarity and signal variability from mobile operation. This demands robust algorithms to identify neural signatures amid environmental noise to enhance decoding accuracy. 
\textbf{Self-supervised learning} (SSL) provides a principled framework for learning noise-invariant representations directly from unlabeled EEG. By designing pretext tasks that exploit inherent electrophysiological pattern~\cite{weng_self-supervised_2025}, SSL models can discover neural signatures resilient to environmental interference. However, traditional voltage reconstruction objective has been reported to underperform due to inherent signal instability~\cite{jiang2024large}. We propose that reconstructing temporally aggregated (average-pooled) representations is more physiologically meaningful. Specifically, by forcing the model to reconstruct the average-pooled temporal segments, we argue that the model can integrate neuroinformative insights while attenuating transient noise, thus capturing sustained neural processes underlying cognitive states. For downstream classification, we integrate \textbf{supervised learning} (SL) to complement SSL's physiological grounding. While SSL reveals fundamental neural dynamics, SL optimizes discriminative boundaries by amplifying task-specific domain-relevant representations. The SSL-SL synergy forms a \textbf{computational duality}: SSL's bottom-up feature extraction and SL's top-down decision refinement jointly emulate the human neurocognitive processing stream, where sensory-level representations are progressively modulated by task-specific demands.

The main contribution of this paper is summarized below:

\textbullet We propose a novel SSL reconstruction objective based on average-pooled EEG, adapting to EEG's intrinsic non-stationary nature.

\textbullet We implement the first joint SSL-SL framework deployed on wearable mobile EEG headsets (Muse), establishing a portable paradigm for ecological cognitive load monitoring.

\section{Methodology} 

\textbf{Encoder}: We deploy parallel convolutions with increasing kernel lengths $(1\times16)$, $(1\times32)$, $(1\times64)$, $(1\times128)$ to capture spectral features $>$16Hz, 8Hz, 4Hz, and 2Hz (since sampling rate is 256Hz). Concatenated features undergo batch normalization and ELU activation, followed by dual temporal pooling (variance + average)~\cite{ma2024attention}. Pooled streams merge via $(2\times1)$ convolution to form latent representations.

\textbf{Dual-Path Decoding}: 
- \textit{SSL Path}: Reconstructs average-pooled signals via two linear layers
- \textit{SL Path}: Projects latent features to logits via 1D convolution

\textbf{Losses}: 
Losses are tailored individually for SSL and SL components. The \textbf{self-supervised} reconstruction loss $\mathcal{L}_{\text{pool}}$ measures the difference between true average pooled EEG $\bm{X}_{\text{avg\_pooled}}$ and reconstructed average pooled EEG $\hat{\bm{X}}_{\text{avg\_pooled}}$ to force model to learn temporally robust features, easing the adversial effect of non-stationarity in EEG. For a given EEG segment, the average-pooled loss is defined as below:
\begin{equation}
\mathcal{L}_{\text{pool}} = \frac{1}{C \times T'} \sum_{c=1}^{C} \sum_{t=1}^{T'} \left( X_{\text{avg\_pooled}}^{(c,t)} - \hat{X}_{\text{avg\_pooled}}^{(c,t)} \right)^2
\end{equation}
where $C$ is the number of channels, $T'$ is the pooled dimension of EEG signals.

The \textbf{supervised} cross-entropy loss refines the features into task-specific representations. For any given EEG segment, this is defined by:
\begin{equation}
\mathcal{L}_{\text{CE}} = -\sum_{k=1}^{K} y_{k} \log\left( p_{k} \right)
\end{equation}
where $K$ indicates the number of possible categories (three in this case, namely low, medium or high level of cognitive load), and $p$ is computed via softmax normalization of the SL decoded logits.

The \textbf{overall loss} for a given sample is calculated as follows:
\begin{equation}
\mathcal{L} =  \mathcal{L}_{\text{CE}} + \lambda \mathcal{L}_{\text{pool}} 
\end{equation}
where $\lambda$ is the SSL loss strength, which is set to 0.5 empirically.

\section{Experiment and Analysis} 

\subsection{Dataset and Experimental Setup}
We utilize the CL-Drive dataset~\cite{muse_cldrive} captured using Muse S headbands during driving simulations. Reported cognitive load scores from 21 participants are categorized into 3 levels (low/medium/high) for each 10-second intervals. To maximize label-temporal relevance, models received the final 2-second segment of each interval. 

EEG data is bandpass-filtered (0.1-75 Hz) with 60Hz notch filtering. Training employed AdamW (lr=0.001) with 100 epochs and batch size 64. Evaluation follows leave-one-subject-out (LOSO) protocol: each participant served as test set once, mimicking real-world new-user scenarios. Reported metrics are averaged performance across all test subjects.

\begin{table}[htbp]
  \centering
  \caption{Comparison of LOSO Performance across Models}
  \label{tab:model_comparison}
  \begin{tabular}{
    l              
    S[table-format=2.2] 
    S[table-format=2.2]
    S[table-format=2.2]
    S[table-format=2.2]
  }
    \toprule
    \textbf{Models}    & \textbf{EEGNet} & \textbf{Deep ConvNet} & \textbf{ResNet} & \textbf{MuseCogNet} \\
    \midrule
    Accuracy (\%) & 57.07 & 57.07 & 58.41 & \textbf{62.68} \\
    F1 score (\%) & 52.09 & 51.25 & 51.61 & \textbf{57.24} \\
    \bottomrule
  \end{tabular}
  \vspace{0.5em} 

\end{table}

\begin{table}[htbp]
  \centering
  \caption{Ablation Results}
  \label{tab:model_ablation}
  \begin{tabular}{
    l              
    S[table-format=2.2] 
    S[table-format=2.2]
    S[table-format=2.2]
    S[table-format=2.2]
  }
    \toprule
    \textbf{Models}     & \textbf{W/O SSL} & \textbf{MuseCogNet}\\
    \midrule
    Accuracy (\%) & 60.77 & \textbf{62.68}\\
    F1 score (\%) & 54.60 & \textbf{57.24}\\
    \bottomrule
  \end{tabular}
  \vspace{0.5em} 

\end{table}

\subsection{Analysis and Discussion}
The comparative analysis reveals distinct performance advantages of our proposed MuseCogNet. As seen in Table~\ref{tab:model_comparison}, established models like EEGNet~\cite{lawhernEEGNetCompactConvolutional2018} and Deep ConvNet~\cite{schirrmeisterDeepLearningConvolutional2018} demonstrate competent baseline performance (Accuracy: 57.07\%, F1: 51.25-52.09\%), validating their status as canonical architectures for EEG decoding. Meanwhile, our implementation of ResNet achieves 58.41\% accuracy, closely matching the original CL-Drive benchmark(58.13\%)~\cite{muse_cldrive}, confirming proper experimental replication. Critically, MuseCogNet achieves superior performance of 62.68\% accuracy and 57.24\% F1, representing \textbf{absolute gains of 4.27\% accuracy and 5.63\% F1} over the strongest baseline (ResNet). We attribute this significant margin to our novel joint learning paradigm enhanced by average-pooled reconstruction objective, where the SSL component learns steady neurodynamic patterns that conventional supervised CNNs fail to capture. The ablation study (Table~\ref{tab:model_ablation}) provides definitive evidence: Removing SSL components causes \textbf{performance degradation of 1.91\% accuracy and 2.64\% F1}, confirming that our average-pooled reconstruction objective contributes substantially to feature robustness. 

Overall, MuseCogNet enables reliable cognitive load monitoring on portable EEG by leveraging SSL-SL synergy—where SSL's bottom-up feature extraction and SL's top-down discriminant refinement exhibit functional similarities to cortical sensory-cognitive processing hierarchies. This joint processing architecture not only achieves superior performance but also suggests intriguing parallels to human attentional mechanisms worthy of future exploration.

\bibliographystyle{IEEEtran}
\bibliography{IEEEabrv,citations}

\end{document}